Use the exact values you enter in previous answer(
Answer #1

10 20 40 50 60 70 90 100 0

object distance p = 81 cm

radius of curvature R = 2 cm

image distacne q = ?

from mirror equation


1/p + 1/q = 2/R

1/81 + 1/q = 2/27


q = 16.2 cm


(c)


image is real


(d)

hi/ho = q/p

hi/9 = 16.2/81

height = 1.8 cm <<<<<<==ANSWER